# 2D Gaussian Splatting for Geometrically Accurate Radiance Fields

### v2

@@ 4.2. Splatting
One common strategy for rendering 2D Gaussians is to project the 2D Gaussian primitives onto the image space using the affine approximation of the perspective projection (Zwicker et al., 2001b, a). However, as noted in  (Zwicker et al., 2004), this projection is only accurate at the center of the Gaussian and has increasing approximation error with increased distance to the center. To address this issue, Zwicker et al. proposed a formulation based on homogeneous coordinates. Specifically, projecting the 2D splat onto an image plane can be described by a general 2D-to-2D mapping in homogeneous coordinates. Let $\mathbf{W}\in{4\times 4}$ be the transformation matrix from world space to screen space. The screen space points are hence obtained by
(7) $\mathbf{x}=(xz,yz,z,1)^{\mathrm{T}}=\mathbf{W}P(u,v)=\mathbf{W}\mathbf{H}(u,v,1,1)^{\mathrm{T}}$
where $\mathbf{x}$ represents a homogeneous ray emitted from the camera and passing through pixel $(x,y)$ and intersecting the splat at depth $z$. To rasterize a 2D Gaussian, Zwicker et al. proposed to project its conic into the screen space with an implicit method using $\mathbf{M}=(\mathbf{W}\mathbf{H})^{-1}$. However, the inverse transformation introduces numerical instability especially when the splat degenerates into a line segment (i.e., if it is viewed from the side). To address this issue, previous surface splatting rendering methods discard such ill-conditioned transformations using a pre-defined threshold. However, such a scheme poses challenges within a differentiable rendering framework, as thresholding can lead to unstable optimization. To address this problem, we utilize an explicit ray-splat intersection inspired by  (Sigg et al., 2006).
@@
We efficiently locate the ray-splat intersections by finding the intersection of three non-parallel planes, a method initially designed for specialized hardware (Weyrich et al., 2007). Given an image coordinate $\mathbf{x}=(x,y)$, we parameterize the ray of a pixel as the intersection of two orthogonal planes: the x-plane and the y-plane. Specifically, the x-plane is defined by a normal vector $(-1,0,0)$ and an offset $x$. Therefore, the x-plane can be represented as a 4D homogeneous plane $\mathbf{h}_{x}=(-1,0,0,x)^{\mathrm{T}}$. Similarly, the y-plane is $\mathbf{h}_{y}=(0,-1,0,y)^{\mathrm{T}}$. Thus, the ray $\mathbf{x}=(x,y)$ is determined by the intersection of the $x$-plane and the $y$-planes.
Next, we transform both planes into the local coordinates of the 2D Gaussian primitives, the $uv$-coordinate system. Note that transforming points on a plane using a transformation matrix $\mathbf{M}$ is equivalent to transforming the homogeneous plane parameters using the inverse transpose $\mathbf{M}^{-\mathrm{T}}$. Therefore, applying $\mathbf{M}=(\mathbf{W}\mathbf{H})^{-1}$ is equivalent to $(\mathbf{W}\mathbf{H})^{\mathrm{T}}$, eliminating explicit matrix inversion and yielding:
@@
As introduced in  Section 4.1, points on the 2D Gaussian plane are represented as $(u,v,1,1)$. At the same time, the intersection point should fall on the transformed $x$-plane and $y$-plane. Thus,
@@
where $\mathbf{h}_{u}^{i},\mathbf{h}_{v}^{i}$ are the $i$-th parameter of the 4D homogeneous plane parameters. We obtain the depth $z$ of the intersected points via Eq. 7 and evaluate the Gaussian value with Eq 6.
@@
where $\mathbf{u}(\mathbf{x})$ is given by (10) and $\mathbf{c}$ is the projection of center $\mathbf{p}_{k}$. Intuitively, $\hat{\mathcal{G}}(\mathbf{x})$ is lower-bounded by a fixed screen-space Gaussian low-pass filter with center $\mathbf{c}$ and radius $\sigma$. In our experiments, we set $\sigma=\sqrt{2}/2$ to ensure sufficient pixels are used during rendering.

### v3

@@ 4.2. Splatting
One common strategy for rendering 2D Gaussians is to project the 2D Gaussian primitives onto the image space using the affine approximation of the perspective projection (Zwicker et al., 2001b, a). However, as noted in  (Zwicker et al., 2004), this projection is only accurate at the center of the Gaussian and has increasing approximation error with increased distance to the center. To address this issue, Zwicker et al. proposed a formulation based on homogeneous coordinates. Specifically, projecting the 2D splat onto an image plane can be described by a general 2D-to-2D mapping in homogeneous coordinates. Let $\mathbf{W}\in{4\times 4}$ be the combined transformation matrix from world space to screen space. The screen space points are hence obtained by
(7) $\mathbf{x}=(xz,yz,z,z)^{\mathrm{T}}=\mathbf{W}P(u,v)=\mathbf{W}\mathbf{H}(u,v,1,1)^{\mathrm{T}}$
where $\mathbf{x}$ represents a homogeneous ray emitted from the camera and passing through pixel $(x,y)$ and intersecting the splat at depth $z$. To rasterize a 2D Gaussian, Zwicker et al. proposed to project its conic into the screen space with an implicit method using $\mathbf{M}=(\mathbf{W}\mathbf{H})^{-1}$. However, the inverse transformation introduces numerical instability, especially when the splat degenerates into a line segment (i.e., if it is viewed from the side). To address this issue, previous surface splatting rendering methods discard such ill-conditioned transformations using a predefined threshold (Zwicker et al., 2004). However, such a scheme poses challenges within a differentiable rendering framework, as thresholding can lead to unstable optimization. To address this problem, we utilize an explicit ray-splat intersection inspired by  (Sigg et al., 2006).
@@
We efficiently locate the ray-splat intersections by finding the intersection of three non-parallel planes, a method originally designed for specialized hardware (Weyrich et al., 2007). Given an image coordinate $\mathbf{x}=(x,y)$, we parameterize the ray of a pixel in the projective space as the intersection of two orthogonal planes: the x-plane and the y-plane. Specifically, the x-plane is defined by a normal vector $(-1,0,0)$ and an offset $x$. The x-plane can be represented as a 4D homogeneous plane $\mathbf{h}_{x}=(-1,0,0,x)^{\mathrm{T}}$. Similarly, the y-plane is $\mathbf{h}_{y}=(0,-1,0,y)^{\mathrm{T}}$. Thus, the ray $\mathbf{x}=(x,y)$ is determined by the intersection of the two planes.
Next, we transform both planes into the local coordinates of the 2D Gaussian primitives, the $uv$-coordinate system. Note that transforming points on a plane using a transformation matrix $\mathbf{M}$ is equivalent to transforming homogeneous plane parameters using the inverse transpose $\mathbf{M}^{-\mathrm{T}}$ (Blinn, 1977). Therefore, applying $\mathbf{M}=(\mathbf{W}\mathbf{H})^{-1}$ is equivalent to $(\mathbf{W}\mathbf{H})^{\mathrm{T}}$, eliminating explicit matrix inversion and yielding:
@@
As introduced in  Section 4.1, points on the 2D Gaussian plane are represented as $(u,v,1,1)$. At the same time, the intersection point should fall in the transformed $x$-plane and $y$-plane. Thus,
@@
where $\mathbf{h}_{u}^{i},\mathbf{h}_{v}^{i}$ are the $i$-th parameter of the 4D plane. Note that $\mathbf{h}_{u}^{3}$ and $\mathbf{h}_{v}^{3}$ are always zero according to Eq. 5. Once we obtain the local coordinates $(u,v)$, we can calculate the depth $z$ of the intersected points using Eq. 7 and evaluate the Gaussian value with Eq. 6.
@@
where $\mathbf{u}(\mathbf{x})$ is given by Eq. 10 and $\mathbf{c}$ is the projection of center $\mathbf{p}_{k}$. Intuitively, $\hat{\mathcal{G}}(\mathbf{x})$ is lower-bounded by a fixed screen-space Gaussian low-pass filter with center $\mathbf{c}$ and radius $\sigma$. In our experiments, we set $\sigma=\sqrt{2}/2$ to ensure sufficient pixels are used during rendering.